The heights of female students at a particular University are distributed with mean of 169 cm


and a standard deviation of 9 cm.


(a) Given that 80% of these female students have a height less than b cm, find the of b.


(b) Given than 60% of these female students have height greater than 5 cm, find the value of s.

1
Expert's answer
2022-01-14T04:54:05-0500

"X"~"N(169, 9^2)"

a) "P(X<b)=0.8\implies P(169+9N(0,1)<b)=0.8\implies P(N(0,1)<{\frac {b-169} 9})=0.8\implies {\frac {b-169} 9}=0.84\implies b=176.56"


b) As i understand, there must be s cm not 5 cm, otherwise it doesn't make much sense

"P(X>s)=0.6\implies P(169+9N(0,1)>s)=0.6\implies P(N(0,1)>{\frac {s-169} 9})=0.6\implies P(N(0,1)<{\frac {s-169} 9})=0.4\implies {\frac {s-169} 9}=-0.253\implies s=166.72"


All of the probabilities you can find in the table of standard normal distribution, it is pretty easy to find in the internet